# A non-linear Roth theorem for thick Cantor sets

### v1

@@ 1. Introduction
$\{x,x+P_{1}(t),\dots,x+P_{n}(t)\}\subset E.$
In the continuous setting, Krause [4] showed that if $E\subset\mathbb{R}$ has sufficiently large Hausdorff dimension and $P$ is a polynomial of degree at least $2$ with no constant term, then there exists $t\neq 0$ such that
$\{x,x-t,x-P(t)\}\subset E.$
@@
Let $K\subset\mathbb{R}$ be a Cantor set with thickness $\tau>1$, and let $f:\mathbb{R}\to\mathbb{R}$ be a continuously differentiable function satisfying $f(0)=0$ and
$\max\left(\frac{\tau}{\tau+1},\frac{1}{\tau}\right)<f^{\prime}(0)<\min\left(\tau,1+\frac{1}{\tau}\right).$
Then, there exist $x\in\mathbb{R},t>0$ such that
@@ 3.2. The non-linear case
Let $a,b,c>0$. Let $K$ be a Cantor set of thickness $\tau>1$ with largest gap (a,0)(-a,0). Let (b,a)(-b,-a) be the left bridge, and $(0,c)$ the right bridge. Assume also that the left bridge is at least as long as the right; that is, $b-a\geq c$. Let $g:\mathbb{R}\to\mathbb{R}$ be a differentiable function satisfying $g(0)=0$ and
$\frac{1}{\tau}<g^{\prime}(\xi)<1+\frac{1}{\tau}\hskip 18.06749pt\textup{if}\hskip 18.06749pt0<\xi<c.$
Then, we have
$0<a<g(c)<b.$
@@
By the Mean Value Theorem, since $g(0)=0$, there exists $\xi\in(0,c)$ such that $g(c)=g^{\prime}(\xi)(c)$. Then, to show $0<a<g(c)<b$, it is sufficient to show that $a<g^{\prime}(\xi)c<b$, or
$\frac{a}{c}<g^{\prime}(\xi)<\frac{b}{c}.$
Let $B$ be either the left or right bridge of $G$. Then, $\frac{|B|}{|G|}\geq\tau$, or $|B|\geq\tau|G|$ Substituting our intervals, we have $b-a\geq\tau a$ and $c\geq\tau a$. To show the upper bound,
$g^{\prime}(\xi)$ $\leq 1+\frac{1}{\tau}$
$=1+\frac{a}{\tau a}$
$\leq 1+\frac{a}{b-a}$
$\leq\frac{b-a+a}{b-a}$
Since $b-a\geq c$ by hypothesis, this gives $g^{\prime}(\xi)\leq\frac{b}{c}$. To show the lower bound,
$g^{\prime}(\xi)$ $\geq\frac{1}{\tau}$
$=\frac{a}{\tau a}$
Since $\tau a\leq c$, this gives $g^{\prime}(\xi)\geq\frac{a}{c}$. Thus, if $\frac{1}{\tau}<g^{\prime}(\xi)<1+\frac{1}{\tau}$, then $a<g(c)<b$ as desired. ∎
@@
By Lemma 2.4 we may choose $K^{\prime}\subset K$ with $\tau(K^{\prime})\geq\tau$ and $|\text{conv}\>K^{\prime}|<\delta$. Let $G$ be the largest gap of $K^{\prime}$, and let $B_{L},B_{R}$ be the left and right bridges of $K^{\prime}$. There are two cases to consider, depending which bridge is larger.
First, suppose $|B_{L}|\geq|B_{R}|$. After translation, we may asssume G=(a,0)G=(-a,0) for some $a>0$. Let $B_{L}=[-b,-a]$ and $B_{R}=[0,c]$. Then, $\tau(K^{\prime})\geq\tau>1$ and by (3.2), $K^{\prime}$ satisfies Lemma 3.3. Let $K_{1}=K^{\prime}\cap[-b,-a]$, and $K_{2}=K^{\prime}\cap[0,c]$. Since $f^{\prime}(0)\neq 0$, if $\delta$ is sufficiently small, then $f$ is invertable on a neighborhood of $K^{\prime}$. Let $g=f^{-1}$. Next, we use the gap lemma to show that there exists some $t\in-K_{1}\cap g(K_{2})$. By Lemma 2.3, $\tau(K_{1})\geq\tau(K^{\prime})>1$. By Lemmas 2.2 and 2.3, $\tau(g(K_{2}))>\rho\tau(K_{2})\geq\rho\tau\geq 1$. Finally, by Lemma 3.3, we have $0<a<g(c)<b$. Thus, there exists $t\in-K_{1}\cap g(K_{2})$ by the Newhouse Gap Lemma. This implies $t\in K_{1}\subset K$ and $f(t)\in K_{2}\subset K$, so $\{-t,0,f(t)\}\subset K^{\prime}$ as desired.
Finally, we consider the case $|B_{R}|\geq|B_{L}|$. The set $K^{\prime}$ then fits the case considered above. Running the same argument with $f^{-1}$ in place of $f$, we obtain a configuration $\{-s,0,f^{-1}(s)\}\subset-K^{\prime}$ for some $s>0$. Letting $t=f^{-1}(s)$ we have $\{-f(t),0,t\}\subset-K^{\prime}$, or $\{-t,0,f(t)\}\subset K^{\prime}$ as desired.

### v3

@@ 1. Introduction
(1.2) $\{x,x+P_{1}(t),\dots,x+P_{n}(t)\}\subset E.$
In the continuous setting, Krause [4] showed that if $\{P_{1},\dots,P_{n}\}$ is a collection of polynomials satisfying a certain technical assumption, and $E\subset\mathbb{R}$ has $s$-dimensional Hausdorff content bounded away from zero for some $s$ sufficiently large, then there exists $t\neq 0$ such that (1.2) holds.
@@
Let $K\subset\mathbb{R}$ be a Cantor set with thickness $\tau>1$, and let $f:\mathbb{R}\to\mathbb{R}$ be a continuously differentiable function satisfying $f(0)=0$ and $\frac{1}{\tau}<f^{\prime}(0)<\tau$. Then, there exist $x\in\mathbb{R},t>0$ such that
@@ 3.2. The non-linear case
Let $a,b,c>0$ be numbers satisfying
(3.1) $a<c\leq b-a,$
and let $g$ be a function differentiable on $(0,c)$ which satisfies
(3.2) $\frac{a}{c}<g^{\prime}<\frac{c}{a}.$
Let $K$ be a Cantor setwith largest gap (a,0)(-a,0) and left and right bridges [b,a][-b,-a] and $[0,c]$, respectively. Finally, define
$K_{L}=K\cap[-b,-a]\hskip 18.06749pt\textup{and}\hskip 18.06749ptK_{R}=K\cap[0,c].$
then, the sets $K_{L}$ and $g(K_{R})$ are interleaved; that is, neither is contained in a single gap of the other.
@@
The equation (3.2) together with the assumption $g(0)=0$ implies that $g$ is stricly increasing, so
$\mathrm{conv}\thinspace(g(K_{R}))=[0,g(c)].$
By the Mean Value Theorem, there exists $\xi\in(0,c)$ such that $g(c)=g^{\prime}(\xi)c$. The lower bound in (3.2) then implies $g(c)>a$. If $g(c)<b$ also, this completes the proof, as the intervals $\mathrm{conv}\thinspace(-K_{L})=[a,b]$ and $\mathrm{conv}\thinspace(g(K_{R}))=[0,g(c)]$ have alternating endpoints. If instead $g(c)>b$, then clearly $g(K_{R})$ cannot lie in a single gap of $K_{L}$. It remains to show $K_{L}$ does not lie in a single gap of $g(K_{R})$ either. Since $K_{L}\subset\mathrm{conv}\thinspace(g(K_{R}))$ in this case, we cannot have $K_{L}$ in an unbounded gap of $g(K_{R})$. To show it does not lie in a single bounded gap, first note that any bounded gap of $g(K_{R})$ is of the form $g(G)$, where $G$ is a bounded gap of $K_{R}$. Because $\mathrm{conv}\thinspace(K_{R})$ is the left bridge (with respect to $K$) of the gap (a,0)(-a,0), we have $|G|\leq a$. By the mean value theorem and the upper bounds in (3.2) and (3.1), we have
$|g(G)|<\frac{c}{a}|G|\leq c\leq b-a=|\mathrm{conv}\thinspace(-K_{L})|.$
Therefore, $K_{L}$ does not lie in $g(G)$. ∎
@@
By Lemma 2.4 we may choose $K^{\prime}\subset K$ with $\tau(K^{\prime})\geq\tau$ and $|\text{conv}\>K^{\prime}|<\delta$. Let $G$ be the largest gap of $K^{\prime}$, and let $B_{L},B_{R}$ be the left and right bridges of $G$. There are two cases to consider, depending which bridge is larger.
First, suppose $|B_{L}|\geq|B_{R}|$. After translation, we may asssume G=(a,0)G=(-a,0) for some $a>0$. Let $b,c>0$ be such that $B_{L}=[-b,-a]$ and $B_{R}=[0,c]$. We claim that Lemma 3.3 applies with this choice of parameters. To verify the hypotheses of Lemma 3.3, first note that $|B_{L}|\geq|B_{R}|$ implies $b-a\geq c$, giving the upper bound of (3.1). Since $c=|B_{R}|$ and $a=|G|$, we have $\frac{c}{a}\geq\tau>1$ and $\frac{a}{c}\leq\frac{1}{\tau}<1$, giving both the lower bound of (3.1) as well as both bounds of (3.2). If $K_{L}=K^{\prime}\cap[-b,-a]$ and $K_{R}=K^{\prime}\cap[0,c]$, we conclude from Lemma 3.3 that $K_{L}$ and $f^{-1}(K_{R})$ are interleaved. By Lemma 2.3,
$\tau(K_{L})\geq\tau(K^{\prime})>1.$
By Lemmas 2.2 and 2.3,
$\tau(f^{-1}(K_{R}))>\rho\tau(K_{R})\geq\rho\tau\geq 1.$
It follows from the Newhouse gap lemma (Theorem 2.1) that there exists some $t\in-K_{1}\cap f^{-1}(K_{2})$. This implies $t\in K_{1}\subset K$ and $f(t)\in K_{2}\subset K$, so $\{-t,0,f(t)\}\subset K^{\prime}$ as desired.
Finally, we consider the case $|B_{R}|\geq|B_{L}|$. The set $K^{\prime}$ then fits the case considered above. Running the same argument with $f$ in place of $f^{-1}$, we obtain a configuration $\{-s,0,f^{-1}(s)\}\subset-K^{\prime}$ for some $s>0$. Letting $t=f^{-1}(s)$ we have $\{-f(t),0,t\}\subset-K^{\prime}$, or $\{-t,0,f(t)\}\subset K^{\prime}$ as desired.